# Second Robin eigenvalue bounds for Schrödinger operators on Riemannian surfaces

### v1

@@ 1. Introduction
Spectral estimates for operators of the form (1.1) (and of the associated boundary problems below) are closely linked to geometry. A fundamental example is the Jacobi operator governing the second variation of area for minimal and constant mean curvature surfaces, which can be written in Schrödinger form. This connection has long been exploited to derive quantitative constraints on stability, index, and topology in surface theory; see for instance [36, 18, 12, 11, 34, 7, 6, 11, 15, 16] and the references therein.
@@
Our second main theorem gives topological restrictions for compact two-sided free boundary minimal surfaces of low Morse index inside geodesic balls of negatively curved pinched Cartan–Hadamard $3$-manifolds. This establishes a hyperbolic analogue of the index–one topological bound proved by Chen–Fraser–Pang in ambients with $\operatorname{Ric}\geq 0$ and weakly convex boundary [9]. Such restrictions are part of a broader circle of ideas relating index bounds, curvature pinching, and topology for minimal/CMC surfaces; see for example [35, 30, 2, 8, 28] and the references therein.
@@
We also establish upper bounds for first eigenvalues in the closed and Robin settings, including rigidity in the curvature-corrected case (compare with intrinsic counterparts of results in immersed surface theory such as [1, 4]). Finally, we obtain Steklov-type estimates in a coercive regime where the Dirichlet problem is well posed for all boundary data and the Dirichlet-to-Neumann operator is defined on the full trace space; cf. related spectral-geometric frameworks in [14, 37, 20, 10].
@@ 2. Closed surfaces: a first-eigenvalue bound
These estimates may be seen as intrinsic counterparts of results proved in the immersed setting, compare [1, Theorem 2.1] and [4, Theorem 1.1].
@@ 3.1. The first Robin eigenvalue
If equality holds, then $u\equiv 1$ is a first eigenfunction (cf. [5, Theorem 2.4]), hence $L1+\mu_{0}(L)=0$ and $B1=0$, which imply $V$ and $W$ are constant; if $a\neq 0$ this forces $K$ and $\kappa_{g}$ to be constant as well. ∎
@@ 4. Application to free boundary minimal surfaces
Finally, apply the linear isoperimetric inequality of Lee–Seo [25, Theorem 2.10(ii)] (with $n=2$, $k=\sqrt{\kappa_{2}}$, and $H=0$):
@@ 5.2. Variational characterizations
Without assuming $\lambda_{0}^{D}(L_{0})>0$, one can still define a Dirichlet-to-Neumann map for $L_{0}$ on a suitable closed subspace of $L^{2}(\partial\Sigma)$ (and hence a corresponding Steklov spectrum), exactly as in Tran [37]. We restrict to the coercive regime to avoid introducing these domain restrictions.

### v2

@@ 1. Introduction
Spectral estimates for operators of the form (1.1) (and of the associated boundary problems below) are closely linked to geometry. A fundamental example is the Jacobi operator governing the second variation of area for minimal and constant mean curvature surfaces, which can be written in Schrödinger form. This connection has long been exploited to derive quantitative constraints on stability, index, and topology in surface theory; see for instance [7, 8, 12, 13, 16, 17, 19, 36, 38] and the references therein.
@@
Our second main theorem gives topological restrictions for compact two-sided free boundary minimal surfaces of low Morse index inside geodesic balls of negatively curved pinched Cartan–Hadamard $3$-manifolds. This establishes a hyperbolic analogue of the index–one topological bound proved by Chen–Fraser–Pang in ambients with $\operatorname{Ric}\geq 0$ and weakly convex boundary [10] (see also [29]). Such restrictions are part of a broader circle of ideas relating index bounds, curvature pinching, and topology for minimal/CMC surfaces; see for example [2, 9, 30, 32, 37] and the references therein.
@@
We also establish upper bounds for first eigenvalues in the closed and Robin settings, including rigidity in the curvature-corrected case (compare with intrinsic counterparts of results in immersed surface theory such as [1, 5]). Finally, we obtain Steklov-type estimates in a coercive regime where the Dirichlet problem is well posed for all boundary data and the Dirichlet-to-Neumann operator is defined on the full trace space; cf. related spectral-geometric frameworks in [3, 11, 15, 21, 39, 41].
@@ 2. Closed surfaces: a first-eigenvalue bound
These estimates may be seen as intrinsic counterparts of results proved in the immersed setting, compare [1, Theorem 2.1] and [5, Theorem 1.1].
@@ 3.1. The first Robin eigenvalue
If equality holds, then $u\equiv 1$ is a first eigenfunction (cf. [6, Theorem 2.4]), hence $L1+\mu_{0}(L)=0$ and $B1=0$, which imply $V$ and $W$ are constant; if $a\neq 0$ this forces $K$ and $\kappa_{g}$ to be constant as well. ∎
@@ 4. Application to free boundary minimal surfaces
Finally, apply the linear isoperimetric inequality of Lee–Seo [26, Theorem 2.10(ii)] (with $n=2$, $k=\sqrt{\kappa_{2}}$, and $H=0$):
@@ 5.2. Variational characterizations
Without assuming $\lambda_{0}^{D}(L_{0})>0$, one can still define a Dirichlet-to-Neumann map for $L_{0}$ on a suitable closed subspace of $L^{2}(\partial\Sigma)$ (and hence a corresponding Steklov spectrum), exactly as in Tran [39]; see also Zhu [41, Section 6] for a closely related construction via a Dirichlet-to-Robin map (and the associated modified Steklov spectrum) in the presence of Dirichlet kernel. We restrict to the coercive regime to avoid introducing these domain restrictions.
@@ Acknowledgments
Data Availability
Data availability is not applicable to this article, as no data sets were generated or analyzed in the course of this research.
Conflict of Interest
The authors declare that they have no conflict of interest related to this article.